cathrinerose23:

solve the equation and check the solution of 8(x +3)-(5x-1)=8 x=? simplify your answer

Shadow:

\[8(x +3)-(5x-1)=8\] Distribution Property \[8x + 24 -5x +1 = 8\] (For the second set of parentheses, note the minus sign before which is basically -1 times 5x -1 which makes 5x negative, and -1 -> positive 1) Now do like terms \[3x + 25 = 8\] Subtraction Property of Equality \[3x = -17\]| Division Property of Equality \[x = - \frac{ 17 }{ 3 }\]